OpenStudy (anonymous):

x2 + 18x + 76 = 0 Which of the following is the first correct step to write the above equation in the form (x - p)2 = q, where p and q are integers?

OpenStudy (dangerousjesse):

Okay, I'm going to change the equation to just\[x^2+18x+76\]

OpenStudy (dangerousjesse):

18x is the middle term, which means that the coefficient we're going to halve is 18.

OpenStudy (dangerousjesse):

\[18\div2=9\]

OpenStudy (dangerousjesse):

\[9^2=?\]

OpenStudy (dangerousjesse):

\[x^2+18x+76=0\]

OpenStudy (dangerousjesse):

\[x^2+18x+76+???=0+???\]

OpenStudy (dangerousjesse):

add five to both sides\[x^2+18x+81\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (dangerousjesse):

No problem.